Which of the following statements are true ?
Give reasons for your answers.
(a) If A belongs to Mn(R), then rank (A)=det(A).
(b) There is one and only one unitary matrix in
Mn(R).
(c) If U and V are subspaces of a vector space
W over R, then U U V is also a subspace of W.
(d) Given a linear transformation T from
R4 to R6, rank (T) + nullity (T) = 6.
(e) The relation R, defined on the set of lines in
R² by `L1 R L2 iff L1 and L2 intersect', is an
equivalence relation.
1
Expert's answer
2020-12-17T18:30:16-0500

a) False. The matrix "\begin{pmatrix}
1&0\\0&1
\end{pmatrix}" is of rank 2 and determinant 1.

b) False. For example, "\begin{pmatrix}
1&0\\0&1
\end{pmatrix}" and "\begin{pmatrix}
0&1\\1&0
\end{pmatrix}" are unitary matrices.

c) False. For example, take "U" to be the "x"-axis and "V" "y"-axis, both subspaces of "\mathbb{R}^2". Their union includes both (1,0) and (0,1), whose sum, (1,1), is not in the union. Hence, the union is not a vector space.

d) False. According to Rank-Nullity theorem, "\text{rank}(T)+\text{nullity}(T)=\dim(\mathbb{R^4})=4\neq6."

e) False. The relation "R" is not transitive. For example, take "L_1,\;L_2\text{ and } L_3" are the lines of equations "y=x,\;y=-x\text{ and }y=x-1" respectively. We have

"L_1\,R \,L_2\text{ and }L_2\, R\, L_3,\text{ but }L_1\not R\, L_3."